Two Spheres of radius 6cms, having charges 10^-8C, -3×10^-8C. Find the potential of 2 spheres and potential at midpoint of centres if the distance between centres is 2m.
1
Expert's answer
2020-09-16T21:09:07-0400

"q_1=10^{-8}C ;q_2=-3\times 10^{-8}C ;r_1=r_2=6\ cm" "=0.06 \ m"

Potential at the surface of first sphere"=\frac{kq_1}{r_1}=\frac{9\times 10^9\times 10^{-8}}{0.06}=1.5\times 10^3 N/m"

Potential at the surface of first sphere"=\frac{kq_2}{r_2}=\frac{9\times 10^9\times (-3)\times 10^{-8}}{0.06}=-4.5\times 10^3 N/m"

Distance between their centers"=\ 2\ m"

So,distance of mid point from center of each sphere"=d_1=d_2=1\ m"

Potential at the mid point due to first sphere"=\frac{kq_1}{d_1}=\frac{9\times 10^9\times 10^{-8}}{1}=90 N/m"

Potential at the mid point due to second sphere"=\frac{kq_2}{d_2}=\frac{9\times 10^9\times (-3)\times 10^{-8}}{1}=-270 N/m"

Net potential at the mid point will be "=(90+(-270)) N/m=-180 \ N/m"